OpenStudy (anonymous):

Solve the following system of equations using matrices. x = -2y + 6 2x + 2y = 16

OpenStudy (anonymous):

****This will give you half credit at least.**** [1 2] [x]=[9] [1 1] [y] [1] A^-1=1/(ad-bc) A^-1=1/-1 [1 -2] [-1 1] A^-1=[-1 2] [1 -1] [-1 2] [1 2] [x]= [-1 2] [9] [1 -1] [1 1] [y] [1 -1] [1] (-1x1)+(2x1)=1 (-1x2)+(2x1)=0 (1x1)+(-1x1)=0 (1x2)+(-1x1)=1 (-1x9)+(2x1)= -7 (1x9)+(1x1)=10 [1 0] [x]= [-7] [0 1] [y] [10] x= -7 y= 10 (-7)-2(10)=-27 Not 9 X Wrong!!! (-7)+(10)=-3 Not 1 X Wrong!!! If I solve not using matrices. x-2y=9 x+y=1 -x -x x-2y=9 y=-x+1 x-2(-x+1)=9 Distribute x+2x-2=9 Combine like terms. 3x-2=9 Subtract 2 on both sides. 3x=7 Divide by 3 on both sides. x=2 1/3 Substitute 7 into x+y=1 so... (7)+y=1 Subtract 7 on both sides. y=-6 (7)-2(-6)=9 7-12= -5 Not 9 X Wrong!! (7)+(-6)=1 7-6=1 Yes 1 Right!!! What is it that I just can't get the question completely right? I tried two methods and can't seem to find my own mistake...

OpenStudy (anonymous):

I have (-2, 10)